What is the peak value of a sinusoidal alternating current of 4.78 r.m.s amperes

What is the r.m.s value of a rectangular voltage wave with an amplitude of 9.87V

What is the average value of sinusoidal alternating current of 31A maximum value?

An alternating current has a periodic time of 0.03seconds. what is its frequency?

(a)An alternating current is represented by i =70.7 sin520t. Determine

Its frequency

The current 0.0015 second after passing through zero and increasing positively

(b) A sinusoidal ac current is given by i = 100 sin∏t. it will achieve a value of 50A after _____ seconds. 


1
Expert's answer
2021-11-07T16:50:25-0500

What is the peak value of a sinusoidal alternating current of 4.78 r.m.s amperes


"I=\frac{I_0}{\sqrt{2}}\to I_0=I\sqrt{2}=6.76\ (A)"


What is the r.m.s value of a rectangular voltage wave with an amplitude of 9.87V


"U(r.m.s.)=9.87\ (V)"


What is the average value of sinusoidal alternating current of 31A maximum value?


"I_{avr}=0"


An alternating current has a periodic time of 0.03 seconds. what is its frequency?


"f=1/T=1/0.03=33.33\ (Hz)"


(a)An alternating current is represented by i =70.7 sin520t. Determine

Its frequency


"\omega=2\pi f\to f=\omega/(2\pi)=520/(2\cdot3.14)=82.76\ (Hz)"


The current 0.0015 second after passing through zero and increasing positively


???


(b) A sinusoidal ac current is given by i = 100 sin∏t. it will achieve a value of 50A after 0.167 seconds.


"50=100\cdot\sin(\pi t)\to 1/2=\sin(\pi t)\to t=1/6=0.167\ (s)"